# Integrated GARCH-GRU in Financial Volatility Forecasting

### v1

@@ Abstract
In this study, we propose a novel integrated Generalized Autoregressive Conditional Heteroskedasticity–Gated Recurrent Unit (GARCH-GRU) model for financial volatility modeling and forecasting. The model embeds the GARCH(1,1) formulation directly into the GRU cell architecture, yielding a unified recurrent unit that jointly captures both traditional econometric properties and complex temporal dynamics. This hybrid structure leverages the strengths of GARCH in modeling key ‘stylized facts’ of financial volatility, such as clustering and persistence, while utilizing the GRU’s capacity to learn nonlinear dependencies from sequential data. Compared to the GARCH-LSTM counterpart, the GARCH-GRU model demonstrates superior computational efficiency, requiring significantly less training time, while maintaining and improving forecasting accuracy. Empirical evaluation across multiple financial datasets confirms the model’s robust outperformance in terms of mean squared error (MSE) and mean absolute error (MAE) relative to a range of benchmarks, including standard neural networks, alternative hybrid architectures, and classical GARCH-type models. As an application, we compute Value-at-Risk (VaR) using the model’s volatility forecasts and observe lower violation ratios, further validating the predictive reliability of the proposed framework in practical risk management settings.
Keywords: Volatility forecasting, Deep Learning, Neural Networks, GARCH models, Gated Recurrent Units, Long Short Term Memory, Value at Risk, Risk Management.
@@ 1 Introduction
Traditional econometric approaches, particularly the Autoregressive Conditional Heteroskedasticity (ARCH) model introduced by Engle (1982) and its generalization, the Generalized ARCH (GARCH) model proposed by Bollerslev (1986), laid the foundation for modern volatility modeling. These models, grounded in strong statistical theory, have proven effective in modeling time-varying conditional variance and capturing volatility clustering. Their interpretability, tractability, and widespread applicability have established them as benchmarks in financial econometrics. However, GARCH-family models rely on linear parametric structures and are often limited in their ability to accommodate non-linear dynamics, structural breaks, and extreme events commonly observed in financial data (Hansen and Lunde (2005)). Several extensions, including asymmetric GARCH variants, long-memory models such as FIGARCH, and more recently, fractional affine multi-component GARCH frameworks, have sought to address some of these limitations. For example, Augustyniak et al. (2022) integrate the Heston-Nandi option pricing framework with FIGARCH dynamics to jointly capture both short- and long-memory components in volatility evolution.
@@
With the advent of high-frequency data and advances in computational techniques, machine learning and deep learning models have emerged as powerful alternatives for volatility forecasting. Models such as recurrent neural networks (RNNs), long short-term memory (LSTM) networks (Hochreiter and Schmidhuber (1997)), and gated recurrent units (GRUs) (Chung et al. (2014)) offer a data-driven approach to uncover complex, nonlinear temporal dependencies in financial time series. These neural architectures have demonstrated superior forecasting accuracy in many settings by flexibly adapting to changing market regimes and capturing intricate dependencies that traditional econometric models may miss. Despite their empirical success, however, deep learning models often suffer from limited interpretability and lack the theoretical grounding inherent in econometric models. Moreover, the use of LSTM-based volatility models in practice is often hampered by computational inefficiency and over-parameterization, especially in latency-sensitive applications such as high-frequency trading.
@@
Recent efforts have attempted to bridge the gap between statistical rigor and data-driven flexibility by combining GARCH-type models with deep learning architectures. Most existing hybrid models have focused on integrating GARCH with LSTM networks. While such hybrids have demonstrated improved predictive performance, the substantial computational overhead associated with LSTMs, which stem from their intricate gating mechanisms and high parameter count, poses practical limitations, particularly in latency-sensitive applications such as high-frequency trading. In contrast, the GRU model offers a more computationally efficient alternative due to its streamlined architecture, while retaining strong capabilities in modeling temporal dependencies (Chung et al. (2014)). Moreover, many existing hybrid approaches adopt a static architecture in which GARCH-generated residuals or conditional variances are simply fed as exogenous inputs to neural networks. Such designs fail to fully exploit the potential for dynamic and reciprocal interactions between econometric structures and learned representations. A notable advancement by Zhao et al. (2024) proposes embedding GARCH regression directly into LSTM cells, effectively replacing the output gate and allowing for deeper integration of volatility dynamics. However, the use of scalar hidden states and the implicit assumption of zero-mean innovations in their model limit its ability to capture external market influences and the complex, high-dimensional temporal dependencies characteristic of modern financial environments.
@@
This study addresses the aforementioned research gaps by introducing a novel integrated GARCH-GRU framework that unifies the interpretability and statistical rigor of traditional econometric models with the nonlinear learning capabilities of deep neural networks. Specifically, we design an architecture that embeds GARCH(1,1) dynamics directly into GRU cells, preserving the gating mechanisms and hidden state transitions intrinsic to the GRU structure. This embedded design enables the model to jointly learn key volatility ‘stylized facts’, such as clustering and persistence, alongside complex temporal patterns in an end-to-end training paradigm, where both the GARCH parameters and GRU weights are optimized simultaneously. The architecture further supports hierarchical feature fusion through stacked integrated layers, facilitating multi-scale representation learning that captures heterogeneous dynamics across time horizons. To evaluate the robustness and generalizability of the proposed framework, we conduct extensive empirical experiments on a diverse set of financial assets, including the S&P 500 Index, the Dow Jones Industrial Average (DJI), and the NASDAQ Composite Index, under varying market regimes and volatility conditions.
The rest of this paper is structured as follows. Section 2 provides a detailed review of related literature in volatility modeling, spanning both econometric and machine learning paradigms. Section 3 presents the methodological framework, including the preliminaries and theoretical formulation of the GARCH-GRU model. Section 4 details the empirical study, describing the datasets, risk metrics, experimental setup, and comparative evaluation against benchmark models. Section 5 concludes the paper by summarizing key findings, discussing limitations, and outlining future directions for extending the proposed framework.
@@ 2 Related Works
Classical econometric models, such as the Autoregressive Conditional Heteroskedasticity (ARCH) model introduced by Engle (1982) and its generalization, the Generalized Autoregressive Conditional Heteroskedasticity (GARCH) model by Bollerslev (1986), laid the groundwork for capturing time-varying volatility and volatility clustering in asset returns. These models are grounded in a well-defined probabilistic framework and have served as the benchmark in both academic and applied settings. To accommodate empirical regularities beyond volatility clustering, numerous extensions to the basic GARCH framework have been introduced. The Integrated GARCH (IGARCH) model by Engle and Bollerslev (1986) incorporates a unit root in the conditional variance process, allowing for shocks to have persistent effects on volatility. The GARCH-in-Mean (GARCH-M) model proposed by Engle et al. (1987) explicitly incorporates conditional variance into the mean equation, capturing the risk-return trade-off. To address the observed asymmetry in volatility responses, where negative returns often lead to larger increases in volatility than positive ones, Nelson (1991) introduced the Exponential GARCH (EGARCH) model, which models the logarithm of the variance and naturally accommodates asymmetric effects. Similarly, the GJR-GARCH model by Glosten et al. (1993) includes an indicator function to differentiate the impact of positive and negative shocks, while the Threshold GARCH (TGARCH) model developed by Zakoian (1994) models the conditional standard deviation and also captures leverage effects. Recognizing the limitations of univariate models in capturing interdependencies across multiple financial time series, multivariate GARCH-type models have also been developed. For instance, the bivariate ARCH model introduced by Engle et al. (1984) represents an early attempt to model co-movements in volatility. The Baba-Engle-Kraft-Kroner (BEKK) model formulated by Engle and Kroner (1995) provides a flexible yet computationally tractable framework for multivariate volatility dynamics. Further extensions include the Varying Correlation MGARCH (VC-MGARCH) model proposed by Tse and Tsui (2002), which allows for time-varying conditional correlations, and the regime-switching MGARCH specification of So and Tse (2009), which accounts for nonlinear shifts in volatility regimes and better accommodates extreme events and tail risk behavior in multivariate settings.
@@
In response to these limitations, deep learning approaches have been increasingly applied to volatility modeling and forecasting. Recurrent neural networks (RNNs), especially Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs), have demonstrated strong capabilities in learning long-range dependencies and nonlinear dynamics in time series data. The LSTM model, introduced by Hochreiter and Schmidhuber (1997) and further extended by Gers et al. (2000) and Fischer and Krauss (2018), has shown empirical superiority in capturing sequential patterns in financial datasets. In a comparative study, Liu (2019) found that LSTM-based models outperform traditional GARCH variants for volatility prediction over longer forecast horizons. GRU, empirically validated by Chung et al. (2014), offers a more computationally efficient alternative to LSTM while maintaining comparable performance in sequence modeling tasks. It has been successfully employed in various financial prediction tasks, including the Two-Stream GRU (TGRU) model proposed by Lien Minh et al. (2018), which integrates bidirectional recurrent mechanisms and achieves improved predictive accuracy over classical models.
@@
Despite their predictive power, deep neural networks often lack financial interpretability and theoretical grounding. Their ‘black-box’ nature poses challenges for transparency, risk attribution, and regulatory acceptability. Furthermore, they do not explicitly model conditional heteroskedasticity, and thus may overlook important volatility-specific features that are well captured by econometric models.
@@
To reconcile the statistical interpretability of GARCH with the representational flexibility of deep learning, a growing body of literature has explored hybrid approaches. Early efforts typically relied on sequential or decoupled modeling pipelines, where GARCH models were used to extract residuals or conditional variances, which were then passed as inputs into neural networks for downstream prediction. Notable examples of such decoupled architectures include the hybrid models developed by Roh (2007), Tseng et al. (2008), Hajizadeh et al. (2012), and Kristjanpoller et al. (2014); Kristjanpoller and Minutolo (2016); Kristjanpoller and Hernández (2017), as well as Kim and Won (2018). While effective in improving forecasting accuracy, these approaches typically involve separate training processes for the statistical and neural components, limiting their capacity to learn joint representations and often resulting in inefficiencies and suboptimal integration. Recent studies have proposed end-to-end hybrid architectures that attempt to jointly optimize econometric and deep learning components. For instance, Zhao et al. (2024) introduced an embedded GARCH-LSTM model that integrates the GARCH regression mechanism within LSTM cells by replacing the output gate. This approach represents a significant step toward unified modeling, enabling direct interaction between GARCH dynamics and temporal feature extraction. However, this model imposes strong assumptions, such as scalar hidden state representations and zero-mean innovations, which restrict its expressiveness and ability to model complex market behaviors and external covariate influences.
@@ 3.1.1 GARCH Model
The GARCH model (Bollerslev (1986)) includes lagged conditional variances along with lagged squared residuals, effectively capturing volatility clustering and persistence in financial time series. Assuming a constant mean return and given the information set $\psi_{t-1}$, the model is specified as:
$r_{t}=\mu+\epsilon_{t}$ (1)
$\epsilon_{t}|\psi_{t-1}\sim\mathcal{D}(0,\sigma_{t}^{2})$ (2)
$\sigma_{t}^{2}=\omega_{0}+\sum_{i=1}^{p}\alpha_{i}\epsilon_{t-i}^{2}+\sum_{j=1}^{q}\beta_{j}\sigma^{2}_{t-j}$ (3)
where $\mathcal{D}(0,\sigma_{t}^{2})$ typically denotes a normal or Student’s t-distribution. The model parameters satisfy $\omega>0$, $\alpha_{i}\geq 0$, $\beta_{j}\geq 0$, and $\sum_{i=1}^{p}\alpha+\sum_{j=1}^{q}\beta<1$ to ensure positivity and weak stationarity of the conditional variance process. The coefficients $\alpha_{i}$ capture the short-term impact of shocks, while the coefficients $\beta_{j}$ reflect the persistence of volatility over time.
@@ 3.1.2 Recurrent Neural Networks
where $x_{t}$ is the input vector at time $t$, $h_{t-1}$ is the previous hidden state, $c_{t-1}$ is the previous cell state, and $\sigma(\cdot)$ denotes the sigmoid activation function. The memory cell $c_{t}$ acts as an internal state that accumulates information over time, while $h_{t}$ serves as the output at time $t$. The element-wise product is denoted by $\odot$.
@@
3.2 Integrated GARCH-GRU Framework

### v2

# Unified GARCH-Recurrent Neural Networks in Financial Volatility Forecasting

@@ Abstract
In this study, we develop a unified volatility modeling framework that embeds GARCH dynamics directly within recurrent neural networks. We propose two interpretable hybrid architectures, GARCH-GRU and GARCH-LSTM, that integrate the GARCH(1,1) volatility update into the multiplicative gating structure of GRU and LSTM cells. This unified design preserves economically meaningful GARCH parameters while enabling the networks to learn nonlinear temporal dependencies in financial time series. Comprehensive out-of-sample evaluations across major U.S. equity indices show that both models consistently outperform classical GARCH specifications, pipeline-style hybrids, and neural baselines such as the Transformer across multiple metrics (MSE, MAE, SMAPE, and out-of-sample $R^{2}$). Within this family, the GARCH-GRU achieves the strongest accuracy–efficiency tradeoff, training nearly three times faster than GARCH-LSTM while maintaining comparable or superior forecasting accuracy under normal market conditions and delivering stable and economically plausible parameter estimates. The advantages persist during extreme market turbulence. In the COVID-19 stress period, both architectures retain superior forecasting accuracy and deliver well-calibrated 99% Value-at-Risk forecasts, achieving lower violation ratios and competitive Pinball losses relative to all benchmarks. Overall, the findings underscore the effectiveness of embedding GARCH dynamics within recurrent neural architectures, yielding models that are accurate, efficient, interpretable, and robust for real-world risk-aware volatility forecasting.
Keywords: Finance; Volatility forecasting; Deep Learning; Recurrent Neural Networks; GARCH models.
@@ 1 Introduction
Traditional econometric approaches, particularly the Autoregressive Conditional Heteroskedasticity (ARCH) model introduced by engle1982autoregressive and its generalization, the Generalized ARCH (GARCH) model proposed by bollerslev1986generalized, laid the foundation for modern volatility modeling. These models, grounded in strong statistical theory, have proven effective in modeling time-varying conditional variance and capturing volatility clustering. Their interpretability, tractability, and widespread applicability have established them as benchmarks in financial econometrics. However, GARCH-family models rely on linear parametric structures and are often limited in their ability to accommodate non-linear dynamics, structural breaks, and extreme events commonly observed in financial data (hansen2005forecast). Extensions such as asymmetric GARCH variants, long-memory models such as fractionally integrated GARCH (FIGARCH), and more recently, fractional affine multi-component GARCH frameworks, offer improvements but still rely on fixed functional forms. Recent findings like (augustyniak2022long) suggest that volatility exhibits multi-component persistence and structural heterogeneity that standard GARCH specifications cannot fully capture, underscoring the need for more adaptive and expressive modeling approaches.
@@
Meanwhile, machine learning and deep learning architectures have emerged as powerful alternatives for volatility forecasting. Models such as recurrent neural networks (RNNs), long short-term memory (LSTM) networks (hochreiter1997long), and gated recurrent units (GRUs) (chung2014empirical) can flexibly approximate nonlinear temporal dependencies and adapt to shifting market regimes. Recent empirical evidence further documents that recurrent deep learning architectures can effectively model nonlinear financial time series, including stock-return dynamics and market-level predictive signals (ThomasChristopher2017). Despite these advances, however, deep learning models often suffer from limited interpretability and lack the theoretical grounding inherent in econometric models. Their internal dynamics do not explicitly encode the ‘stylized facts’ that econometricians have documented over decades. Moreover, LSTM-based volatility models can be computationally heavy and over-parameterized, limiting their deployability in latency-sensitive settings.
@@
Recent efforts have attempted to bridge the gap between statistical rigor and data-driven flexibility by combining GARCH-type models with deep learning architectures. Existing work largely focuses on LSTM-based hybrids that provide improved flexibility but still treat GARCH outputs as exogenous features, rather than embedding volatility dynamics directly into the recurrent architecture. This static or feed-forward integration limits the extent to which neural networks can interact with or adaptively modulate econometric components. A recent advancement by zhao2024garch embeds GARCH regression into the LSTM cell, replacing the output gate and enabling deeper integration of volatility signals. However, their formulation uses scalar hidden states and implicitly assumes zero-mean innovations, which restricts the model’s capacity to capture external drivers, multi-dimensional dependence structures, and richer forms of temporal heterogeneity. Beyond structural limitations, LSTM-based hybrids also face practical constraints. The LSTM architecture carries substantial computational overhead due to its intricate gating structure and large parameter count, leading to heavier training costs and slower inference, particularly in latency-sensitive applications such as high-frequency trading. In contrast, the GRU architecture offers a more streamlined gating mechanism with significantly fewer parameters, yielding meaningful improvements in computational efficiency while retaining strong capacity to model nonlinear and temporal dependencies (chung2014empirical). The computational and architectural advantages make GRU-based hybrids particularly appealing for real-world deployment.
@@
In this study, we introduce a unified GARCH–Recurrent Neural Network (GARCH-RNN) framework that integrates the interpretability and statistical rigor of traditional econometric models with the nonlinear learning capabilities of deep neural networks. The proposed framework embeds GARCH(1,1) dynamics directly into neural network (NN) cells, preserving the gating mechanisms and hidden state transitions intrinsic to the NN structure. This embedded design enables the model to jointly learn key volatility ‘stylized facts’, such as clustering and persistence, alongside complex temporal patterns in an unified training paradigm, where both the GARCH parameters and NN weights are optimized simultaneously. The architecture further supports hierarchical feature fusion through stacked integrated layers, facilitating multi-scale representation learning that captures heterogeneous dynamics across time horizons. Within this general framework, we develop two specific model instantiations based on GRU and LSTM architectures. Despite their architectural differences, both variants retain explicit GARCH parameters, which remain economically interpretable even after joint optimization with neural network weights. To evaluate the robustness and generalizability of the proposed framework, we conduct extensive empirical experiments on a diverse set of financial assets, including the S&P 500 Index, the Dow Jones Industrial Average (DJI), and the NASDAQ Composite Index, under varying market regimes and volatility conditions. The results demonstrate that the GARCH-RNN framework substantially improves volatility forecasting accuracy relative to both classical GARCH models and state-of-the-art deep learning baselines. Moreover, the complementary behavior of the GARCH-GRU and GARCH-LSTM variants provides insight into the regime-dependent nature of volatility dynamics and highlights the value of integrating econometric structure with neural sequence models.
The rest of the paper is structured as follows. Section 2 reviews related literature spanning econometric volatility models, deep learning approaches, and recent hybrid methods. Section 3 presents the methodological development of the GARCH-RNN framework, including the preliminaries and the theoretical formulation of the GRU- and LSTM-based variants. Section 4 describes the empirical design, datasets, evaluation metrics, and benchmark models, followed by a comprehensive presentation of the evaluation results. Section 5 concludes with a discussion of the main findings, practical implications, limitations, and avenues for future research.
@@ 2 Related Works
Classical econometric models, such as the ARCH and GARCH models, laid the groundwork for capturing time-varying volatility and volatility clustering in asset returns. These models are grounded in a well-defined probabilistic framework and have served as the benchmark in both academic and applied settings. To accommodate empirical regularities beyond volatility clustering, numerous extensions to the basic GARCH framework have been introduced. The Integrated GARCH (IGARCH) model by engle1986modelling incorporates a unit root in the conditional variance process, allowing for shocks to have persistent effects on volatility. The GARCH-in-Mean (GARCH-M) model proposed by engle1987estimating explicitly incorporates conditional variance into the mean equation, capturing the risk-return trade-off. To address the observed asymmetry in volatility responses, where negative returns often lead to larger increases in volatility than positive ones, nelson1991conditional introduced the Exponential GARCH (EGARCH) model, which models the logarithm of the variance and naturally accommodates asymmetric effects. Similarly, the GJR-GARCH model by glosten1993relation includes an indicator function to differentiate the impact of positive and negative shocks, while the Threshold GARCH (TGARCH) model developed by zakoian1994threshold models the conditional standard deviation and also captures leverage effects. Recognizing the limitations of univariate models in capturing interdependencies across multiple financial time series, multivariate GARCH-type models have also been developed. For instance, the bivariate ARCH model introduced by Engle1984 represents an early attempt to model co-movements in volatility. The Baba-Engle-Kraft-Kroner (BEKK) model formulated by engle1995multivariate provides a flexible yet computationally tractable framework for multivariate volatility dynamics that has been influential in empirical finance application, such as bollerslev2008glossary. Subsequent extensions further enhance the ability to model dynamic cross-asset interactions. The Varying Correlation MGARCH model of Tse2002 captures time‐varying conditional correlations through a parsimonious updating mechanism, while regime-switching MGARCH frameworks such as So2009 incorporate nonlinear transitions across volatility regimes, thereby accommodating the extreme events, structural breaks, and tail-risk behavior widely documented in empirical studies (e.g., poon2003forecasting; andersen2003modeling).
@@
Beyond volatility modeling itself, machine learning methods have shown strong predictive performance across a wide range of financial forecasting tasks. These include equity return prediction using ensemble learners and deep neural networks (krauss2017deep), as well as large-scale risk forecasting problems characterized by nonlinear interactions and high-dimensional covariates (huck2019large). Hybrid neural-network architectures have also been explored in financial contexts. For example, sermpinis2013forecasting develop an adaptive RBF–PSO model that outperforms ARMA and several neural-network benchmarks in exchange-rate forecasting, illustrating how nonlinear and optimization-enhanced architectures can extract complex temporal patterns from financial time series. Building on these advances, deep learning approaches have been increasingly applied to volatility modeling and forecasting. Recurrent neural networks (RNNs), especially Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs), are well suited for capturing long-range dependencies and nonlinear dynamics in time series data. The LSTM model, introduced by hochreiter1997long and further extended by Gers2000 and ThomasChristopher2017, has shown empirical superiority in capturing sequential patterns in financial datasets. In a comparative study, liu2019novel found that LSTM-based models outperform traditional GARCH variants for volatility prediction over longer forecast horizons. GRU, empirically validated by chung2014empirical, offers a more computationally efficient alternative to LSTM while maintaining comparable performance in sequence modeling tasks. It has been successfully employed in various financial prediction tasks, including the Two-Stream GRU (TGRU) model proposed by 8456512, which integrates bi-directional recurrent mechanisms and achieves improved predictive accuracy over classical models.
@@
Despite their remarkable predictive power, deep neural networks often lack financial interpretability and theoretical grounding. Similar concerns have been noted in broader areas of financial risk modeling, where deep learning models, such as those used in bankruptcy prediction, also face persistent challenges related to transparency and explainability (mai2019deep). The inherent ‘black-box’ nature of these architectures poses challenges for transparency, risk attribution, and regulatory acceptability. Furthermore, they do not explicitly model conditional heteroskedasticity, and thus may overlook important volatility-specific features that are well captured by econometric models.
@@
To reconcile the statistical interpretability of GARCH with the representational flexibility of deep learning, a growing body of literature has explored hybrid approaches. Early efforts typically relied on sequential or decoupled modeling pipelines, where GARCH models were used to extract residuals or conditional variances, which were then passed as inputs into neural networks for downstream prediction. Notable examples of such decoupled architectures include the hybrid models developed by roh2007forecasting, tseng2008artificial, hajizadeh2012hybrid, and kristjanpoller2014volatility; kristjanpoller2016forecasting; kristjanpoller2017volatility, as well as kim2018forecasting. While effective in improving forecasting accuracy, these approaches typically involve separate training processes for the statistical and neural components, limiting their capacity to learn joint representations and often resulting in inefficiencies and suboptimal integration. Recent studies have proposed end-to-end hybrid architectures that attempt to jointly optimize econometric and deep learning components. For instance, zhao2024garch introduced an embedded GARCH-LSTM model that integrates the GARCH regression mechanism within LSTM cells by replacing the output gate. This approach represents a significant step towards unified modeling, enabling direct interaction between GARCH dynamics and temporal feature extraction. However, this model imposes strong assumptions, such as scalar hidden state representations and zero-mean innovations, which restrict its expressiveness and ability to model complex market behaviors and external covariate influences.
@@ 3.1.1 GARCH Model
The GARCH model (bollerslev1986generalized) includes lagged conditional variances along with lagged squared residuals, effectively capturing volatility clustering and persistence in financial time series. Assuming a constant mean return and given the information set $\psi_{t-1}$, the model is specified as:
$r_{t}=\mu+\epsilon_{t},$ (1)
$\epsilon_{t}|\psi_{t-1}\sim\mathcal{D}(0,\sigma_{t}^{2}),$ (2)
$\sigma_{t}^{2}=\omega_{0}+\sum_{i=1}^{p}\alpha_{i}\epsilon_{t-i}^{2}+\sum_{j=1}^{q}\beta_{j}\sigma^{2}_{t-j},$ (3)
where $\mathcal{D}(0,\sigma_{t}^{2})$ typically denotes a normal or Student’s t-distribution. The model parameters satisfy $\omega>0$, $\alpha_{i}\geq 0$, $\beta_{j}\geq 0$, and $\sum_{i=1}^{p}\alpha_{i}+\sum_{j=1}^{q}\beta_{j}<1$ to ensure positivity and weak stationarity of the conditional variance process. The coefficients $\alpha_{i}$ capture the short-term impact of shocks, while the coefficients $\beta_{j}$ reflect the persistence of volatility over time.
@@ 3.1.2 Recurrent Neural Networks
where $x_{t}$ is the input vector at time $t$, $h_{t-1}$ is the previous hidden state, $c_{t-1}$ is the previous cell state, and $\sigma(\cdot)$ denotes the sigmoid activation function. The memory cell $c_{t}$ acts as an internal state that accumulates information over time, while $h_{t}$ serves as the output at time $t$. The element-wise product operation (Hadamard product) is denoted by $\odot$.
@@
3.2 Unified GARCH-RNN Framework